# bibliography.bib

### v1

@@ 1 Introduction
The method that fulfils all these criteria is presented in theorem 1 and called exact Fourier acceleration (EFA). It is summarised in algorithms 1, 2 and 3. The idea of Fourier acceleration (FA) reaches back to Ref. [PhysRevD.32.2736], while a scale separation of different force contributions has been introduced in Ref. [URBACH200687] and used in combination with FA in Ref. [Scalettar:2203.01291]. The scale separation is required here for the individual treatment of the quadratic or harmonic part of the action. Recently, in Ref. [Ostmeyer:2023azi] we have introduced a first version of EFA. This work generalises EFA to actions with arbitrary quadratic terms, not just those diagonalisable by Fourier transformation. Moreover, the importance of the trajectory length is identified and the optimal value derived.
@@ 2.2 Exact Fourier acceleration and optimal HMC parameters
In many practical applications $M$ is translationally invariant and can be diagonalised and inverted via Fourier transformation. The idea to use $M^{-1}$ in the kinetic term first arose in this context which is why this choice is typically called Fourier acceleration (FA). We will stick to this historically motivated nomenclature, however theorem 1 is more generally applicable and “harmonic” acceleration might be a more appropriate name.
@@
This proof is highly instructive in that it provides us with an analytic exact solution to the harmonic EOM. This solution can be used to augment the numerical integration of the EOM by solving the harmonic (or Fourier) part exactly and only treating the anharmonic forces $-\nabla V(x)$ numerically, see algorithms 2 and 3. As in Ref. [Ostmeyer:2023azi], this method is called exact Fourier acceleration (EFA).
@@
In a direct comparison of classical FA [PhysRevD.32.2736, Scalettar:2203.01291] (i.e. using a numerical solution to the EOM) and EFA, one finds that EFA has a number of advantages and no disadvantages. By far the most important advantage is that it allowed for the choice of the optimal trajectory length in theorem 1. It also comes with a computational speed up since the harmonic part of the trajectory does not have to be divided into discrete steps, each requiring additional FFTs. In practice this speed up is often negligible compared to the cost of calculating the anharmonic potential, typically involving fermionic contributions with determinants or inversions. For practical applications it turns out to be more useful that EFA comes with the smallest possible energy violation and thus highest possible acceptance rate without the need to tune the number of discrete update steps.
@@ 2.3 Implication of sub-optimal parameters
If no FA is used, i.e. instead of the Hamiltonian (7) one chooses $\mathcal{H}=\frac{1}{2}p^{2}+S(x)$, then the integrated autocorrelation time $\tau_{\mathrm{int}}$ of an observable $\mathcal{A}$ will be rescaled by the factor
@@
The result of corollary 1 is also found in [KENNEDY1991118], though a slightly different derivation and a less general setting are used.
@@
Thus, choosing a short trajectory length leads to a linear divergence of the autocorrelation time in $1/T$, while neglecting to use FA results in quadratic divergence in the frequency ratio $\omega_{\mathrm{max}}/\omega_{\mathrm{min}}$ or, equivalently, linear divergence in the condition number of $M$.
@@ 2.4 Deviations for strongly anharmonic potentials
This problem can easily be avoided by choosing a variable random trajectory length. No-U-turn sampling [NUTS-2014] is an example of a very elaborate algorithm that effectively chooses a random $T\in[-\pi,\pi]$ without assuming any prior knowledge about the periodicity. A simple and yet effective alternative method is to choose $T$ uniformly from a fixed interval as in Ref. [cohenstead2024smoqydqmcjl]. The theoretically optimal interval is $T\in\left[\nicefrac{{\pi}}{{2}}-\delta,\nicefrac{{\pi}}{{2}}+\delta\right]$ with $0\leq\delta\leq\nicefrac{{\pi}}{{2}}$. For small anharmonic potentials $V(x)$ one can safely choose $\delta=0$ and only for very large $V(x)$ some $\delta>0$ is required. The precise value has to be adjusted during numerical experiments and depends on the specifics of the system. In all simulations presented in this work, a constant trajectory length was found sufficient to obtain unbiased results and this direction has not been further explored.
@@ 3.3 Lattice gauge theory
Our final example is motivated by lattice quantum chromodynamics (QCD), the theory describing quarks and gluons under the strong interaction. FA has been known in simulations of lattice QCD for decades. The original formulation in Refs. [DUANE1986143, DUANE1988101] takes the structure of the underlying gauge group into account explicitly. It cannot be derived from theorem 1 and comes at the cost of numerical inversions or solves. In the following, an alternative idea how theorem 1 can be used in lattice QCD and gauge theories in general will be sketched. This approximate FA does not require numerical solves, but it does not mirror the exact group structure either. The approach is a generalisation of the FA used in Ref. [Borsanyi:2015] for quantum electrodynamics (QED) beyond the abelian U$(1)$ gauge group.
@@
is defined in the usual way as a function of the gauge links $U_{\mu}(n)$. Clearly, the action is not quadratic in the links, so theorem 1 is not applicable without further manipulations. Note that the gauge action is quadratic in the plaquette $S_{G}\sim\sum_{n,\mu<\nu}\operatorname{Tr}\left[(1-U_{\mu\nu}(n))^{\dagger}(1-U_{\mu\nu}(n))\right]$ and this property can be exploited for highly efficient sampling, for instance with the heat bath algorithm [PhysRevD.21.2308]. Unfortunately this ansatz does not generalise to HMC simulations because the EOM for plaquettes and their canonical momenta would inevitably violate unitarity (see section C.1 for more details).
@@
with the lattice momenta $k$. The harmonic gauge matrix $M_{G}$ is defined by its action $M^{\mu\nu}_{G}(k)$ on the fields $y$ after Fourier transformation. On a hyper-cubic lattice in $d$ dimensions every site has $d$ associated links, so the remaining non-diagonal blocks are of dimensionality $d\times d$ and very fast to diagonalise numerically. At this point theorem 1 is directly applicable and momenta can be sampled accordingly. However, the EOM should be integrated numerically by taking the exact links $U_{\mu}(n)$ into account rather than their harmonic approximations, therefore no EFA is possible. An alternative approach to FA in gauge theories is sketched in section C.4.
@@
This band captures the remaining local gauge degrees of freedom. Removing the dynamics of these zero modes, leads to a very natural gauge fixing which has been used for FA in this work (correctness is demonstrated in section C.3). An alternative approach could be to regulate the zeros by adding a small constant to $M_{G}$ the way classical FA typically proceeds.
11111010101010010010010020202020404040406060606080808080100100100100120120120120140140140140160160160160180180180180200200200200Refer to caption
@@
$L$FA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=$0.6$$
1111101010103333666630303030111110101010Refer to caption
@@
$\beta$FA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=\nicefrac{{1}}{{\sqrt{\beta}}}$no FA, $T=1$
@@ Acknowledgements
We thank Evan Berkowitz, Benjamin Cohen-Stead, Stefan Krieg, Tom Luu, and Carsten Urbach for their helpful comments. Marcel Rodekamp deserves special thanks for the very enjoyable and productive joint literature research of pure gauge theory and topology. This work was funded in part by the STFC Consolidated Grant ST/T000988/1 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the CRC 1639/1 NuMeriQS – 511713970. Numerical simulations were undertaken on Barkla (though simulations with EFA could have easily been run on a laptop), part of the High Performance Computing facilities at the University of Liverpool, UK.
@@ C.1 Why heat bath cannot be generalised
The heat bath algorithm [PhysRevD.21.2308, Gattringer:2010zz] is an efficient way to sample new link variables locally in pure gauge theory. It effectively removes the potential rejection in standard Metropolis Hastings sampling. The heat bath uses the property that the gauge action $S_{G}$ is quadratic in the plaquette $U_{\mu\nu}(n)$
@@ C.3 Numerical verification of the approximate FA
As expected, we find excellent agreement between all the methods for the strong coupling regime $\beta\lesssim 4$ and the two numerical methods agree within errors for the entire tested parameter range.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111002222444466668888101010101212121214141414Refer to caption
@@
$\beta$FA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=\nicefrac{{1}}{{\sqrt{\beta}}}$strong coupling expansion
$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$00111122223333444455556666Refer to caption
@@
$\beta$FA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=\nicefrac{{1}}{{\sqrt{\beta}}}$
Figure 5: Plaquette results in SU$(3)$ pure gauge theory on a $10^{4}$ lattice from HMC simulations checked against the strong coupling expansion [PhysRevD.11.2104, PhysRevD.19.2514] using eq. (69). The 16th order expansion was converted to Padé approximation form in order to maximise the range of validity. Left: plaquette expectation value $\left\langle P(\beta)\right\rangle$; Right: absolute value of the deviation of $\left\langle P(\beta)\right\rangle$ from eq. (69).
@@ C.4 Alternative approach sampling the plaquette
From section C.1 we learn that we have to sample in the algebra, as is done in classical lattice QCD HMC simulations for exactly this reason. One alternative option to sampling the links as in section C.2 is to still sample the plaquette like heat bath does. This means that we would also have to choose a subset of links to update at a time, but this subset can be extensive in the volume, so an update would still be global. Expand the plaquette $U_{\mu\nu}(n)=\exp\left(\operatorname{i}X_{n,\mu\nu}\right)$ around unity to obtain $S_{G}\sim\beta X^{2}$. Now sample momenta according to this harmonic part and then evolve the EOM using the force from the full action.

### v2

@@ 1 Introduction
The method that fulfils all these criteria is presented in theorem 1 and called exact Fourier acceleration (EFA). It is summarised in algorithms 1, 2 and 3. The idea of Fourier acceleration (FA) reaches back to Ref. [PhysRevD.32.2736], while a scale separation of different force contributions has been introduced in Ref. [Sexton:1992nu, URBACH200687] and used in combination with FA in Ref. [Scalettar:2203.01291]. The scale separation is required here for the individual treatment of the quadratic or harmonic part of the action. Recently, in Ref. [Ostmeyer:2023azi] we have introduced a first version of EFA. This work generalises EFA to actions with arbitrary quadratic terms, not just those diagonalisable by Fourier transformation. Moreover, the importance of the trajectory length is identified and the optimal value derived.
@@ 2.2 Exact Fourier acceleration and optimal HMC parameters
In many practical applications $M$ is translationally invariant and can be diagonalised and inverted via Fourier transformation. The idea to use $M^{-1}$ in the kinetic term first arose in this context which is why this choice is typically called Fourier acceleration (FA). We will stick to this historically motivated nomenclature, however theorem 1 is more generally applicable and “harmonic” acceleration might be a more appropriate name. The optimal choice of $M^{-1}$ together with $T=\frac{\pi}{2}$ has been used in the special case when $M$ can be inverted with a Fourier transformation [Borsanyi:2015, Cianci:2023vtm]. We also note that the generalised use of $M^{-1}$ has been advocated before [Riemann_Manifold:2011, Neal:2011mrf], but to our knowledge not in conjunction with the optimal trajectory length.
@@
This proof is highly instructive in that it provides us with an analytic exact solution to the harmonic EOM. This solution can be used to augment the numerical integration of the EOM by solving the harmonic (or Fourier) part exactly and only treating the anharmonic forces $-\nabla V(x)$ numerically, see algorithms 2 and 3. As in Ref. [Ostmeyer:2023azi], this method is called exact Fourier acceleration (EFA). Note that the initial formulation of EFA [Ostmeyer:2023azi] did not include the optimal parameter choice as in eqs. 14 and 15. It should not be used since the version introduced in this work is strictly superior. From now on we will use the term EFA exclusively to denote the optimal algorithm introduced here.
@@
In a direct comparison of classical FA [PhysRevD.32.2736, Scalettar:2203.01291, PhysRevD.56.6885, Catterall:2001jg, Flamino:2018jmo, Horsley:2023fhj] (i.e. using a numerical solution to the EOM) and EFA, one finds that EFA has a number of advantages and no disadvantages. By far the most important advantage is that it allowed for the choice of the optimal trajectory length in theorem 1. It also comes with a computational speed up since the harmonic part of the trajectory does not have to be divided into discrete steps, each requiring additional FFTs. In practice this speed up is often negligible compared to the cost of calculating the anharmonic potential, typically involving fermionic contributions with determinants or inversions. For practical applications it turns out to be more useful that EFA comes with the smallest possible energy violation and thus highest possible acceptance rate without the need to tune the number of discrete update steps.
@@ 2.3 Implication of sub-optimal parameters
If no FA is used, i.e. instead of the Hamiltonian (7) one chooses $\mathcal{H}=\frac{1}{2}p^{2}+S(x)$ with fixed trajectory length, then the integrated autocorrelation time $\tau_{\mathrm{int}}$ of an observable $\mathcal{A}$ will be rescaled by the factor
@@
The result of corollary 1 is also found in [KENNEDY1991118], though a slightly different derivation and a less general setting are used. Recently, in Ref. [apers2022hamiltonian] it has been shown that the total computational cost of HMC sampling (up to logarithmic corrections) can be reduced to $\operatorname{\mathcal{O}}\left(\frac{\omega_{\mathrm{max}}}{\omega_{\mathrm{min}}}\right)$ even without FA by ensuring a long enough trajectory length $T_{\text{max}}\gtrsim\omega_{\mathrm{min}}^{-1}$ and sampling $T\in(0,T_{\text{max}}]$ randomly. In this scenario $\tau_{\text{int}}$ remains constant and the cost of a single trajectory scales with $\omega_{\mathrm{max}}/\omega_{\mathrm{min}}$.
@@
Thus, choosing a short trajectory length leads to a linear divergence of the autocorrelation time in $1/T$, while neglecting to use FA at fixed $T$ results in quadratic divergence in the frequency ratio $\omega_{\mathrm{max}}/\omega_{\mathrm{min}}$ or, equivalently, linear divergence in the condition number of $M$. Following Ref. [apers2022hamiltonian], linear divergence of the cost in $\omega_{\mathrm{max}}/\omega_{\mathrm{min}}$ can be achieved without FA using correspondingly long trajectories.
@@ 2.4 Deviations for strongly anharmonic potentials
This problem can easily be avoided by choosing a variable random trajectory length [Mackenzie:1989us]. No-U-turn sampling [NUTS-2014] is an example of a very elaborate algorithm that effectively chooses a random $T\in[-\pi,\pi]$ without assuming any prior knowledge about the periodicity. A simple and yet effective alternative method is to choose $T$ uniformly from a fixed interval as in Ref. [cohenstead2024smoqydqmcjl]. The theoretically optimal interval is $T\in\left[\nicefrac{{\pi}}{{2}}-\delta,\nicefrac{{\pi}}{{2}}+\delta\right]$ with $0\leq\delta\leq\nicefrac{{\pi}}{{2}}$. For small anharmonic potentials $V(x)$ one can safely choose $\delta=0$ and only for very large $V(x)$ some $\delta>0$ is required. The precise value has to be adjusted during numerical experiments and depends on the specifics of the system. In all simulations presented in this work, a constant trajectory length was found sufficient to obtain unbiased results and this direction has not been further explored.
@@
Finally, while theorem 1 guarantees (instant) thermalisation for harmonically dominated actions, no such guarantee is given for general actions. In particular, on non-compact manifolds the HMC (independently of FA) needs to be augmented by a so-called radial update in order to guarantee thermalisation [original_radial_update, Ostmeyer:2024gnh]. Such a radial update introduces non-local steps, e.g. multiplicative updates of the current state, and thus allows to overcome slow diffusive behaviour in remote regions of the phase space. It can also restore ergodicity to the simulations of some actions with phase space separation [Temmen:2024pcm, Temmen_ergodicity].
@@ 3.3 Lattice gauge theory
Our final example is motivated by lattice quantum chromodynamics (QCD), the theory describing quarks and gluons under the strong interaction. FA has been known in simulations of lattice QCD for decades. The original formulation in Refs. [DUANE1986143, DUANE1988101, Davies:1989vh] as well as more recent implementations [Sheta:2021hsd, Huo:2024lns] takes the structure of the underlying gauge group into account explicitly. It cannot be derived from theorem 1 and comes at the cost of numerical inversions or solves. In the following, an alternative idea how theorem 1 can be used in lattice QCD and gauge theories in general will be sketched. This approximate FA does not require numerical solves, but it does not mirror the exact group structure either. The approach is a generalisation of the FA used in Ref. [Borsanyi:2015] for quantum electrodynamics (QED) beyond the abelian U$(1)$ gauge group.
@@
is defined in the usual way as a function of the gauge links $U_{\mu}(n)$. Clearly, the action is not quadratic in the links, so theorem 1 is not applicable without further manipulations. Note that the gauge action is quadratic and diagonal in the plaquette $S_{G}\sim\sum_{n,\mu<\nu}\operatorname{Tr}\left[(1-U_{\mu\nu}(n))^{\dagger}(1-U_{\mu\nu}(n))\right]$ allowing to sample the plaquette directly. This property can be exploited for highly efficient sampling, for instance with the heat bath algorithm [PhysRevD.21.2308]. The idea there is to update some plaquettes while keeping all but one adjacent links fixed. Then the change of the plaquettes directly translates to a change of the variable link. Unfortunately this ansatz does not generalise to HMC simulations because the EOM for plaquettes and their canonical momenta would inevitably violate unitarity (see section C.1 for more details).
@@
with the lattice momenta $k$. The harmonic gauge matrix $M_{G}$ is defined by its action $M^{\mu\nu}_{G}(k)$ on the fields $y$ after Fourier transformation. On a hyper-cubic lattice in $d$ dimensions every site has $d$ associated links, so the remaining non-diagonal blocks are of dimensionality $d\times d$ and very fast to diagonalise numerically. At this point theorem 1 is directly applicable and momenta can be sampled accordingly. However, the EOM should be integrated numerically by taking the exact links $U_{\mu}(n)$ into account rather than their harmonic approximations, therefore no EFA is possible. In summary, the FA algorithm employed here only differs from the well-known standard HMC algorithm [Duane1987, Gattringer:2010zz] in that the kinetic term $\frac{1}{2}p^{\mathsf{T}}\tilde{M}_{G}^{-1}p$ is used instead of $\frac{1}{2}p^{2}$ (with the regularised generalised inverse $\tilde{M}_{G}^{-1}$ defined in section C.2). Correctness of the simulations is demonstrated in section C.3. An alternative approach to FA in gauge theories is sketched in section C.4.
@@
This band captures the remaining local gauge degrees of freedom. Entirely removing the dynamics of these zero modes, leads to a gauge fixing which is potentially error-prone in non-abelian gauge theories [DUANE1988101, Huo:2024lns]. The very simple fix used in this work is to replace the exact zeroes by slow dynamics, see section C.2 for details. An alternative approach could be to regulate the zeros by adding a small constant to $M_{G}$ the way classical FA typically proceeds. Once all momenta can take non-zero values, the simulation has the same ergodicity properties as the classical HMC without FA and can only differ in terms of efficiency.
11111010101010010010010020202020404040406060606080808080100100100100120120120120140140140140160160160160180180180180200200200200Refer to captionFA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=$0.6$$
@@
$L$
1111101010103333666630303030111110101010Refer to captionFA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=\nicefrac{{1}}{{\sqrt{\beta}}}$no FA, $T=1$
@@
$\beta$
@@ 4 Conclusion
Only the combination of all the following points, collected in the EFA approach for the first time, allows for optimal sampling:
  • kinetic ‘mass’ matrix $M^{-1}$ (not only for Fourier invertible cases),
  • HMC trajectory length $T=\frac{\pi}{2}$,
  • exact integration of the harmonic forces, numerical treatment of the rest.
With EFA the HMC becomes equivalent to direct sampling for harmonic actions $S(x)=\frac{1}{2}x^{\mathsf{T}}Mx$ and only perturbatively worse for perturbative anharmonic parts.
@@ Acknowledgements
We thank Evan Berkowitz, Benjamin Cohen-Stead, Sander Gribling, Anthony Kennedy, Stefan Krieg, Tom Luu, and Carsten Urbach for their helpful comments. Marcel Rodekamp deserves special thanks for the very enjoyable and productive joint literature research of pure gauge theory and topology. This work was funded in part by the STFC Consolidated Grant ST/T000988/1 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the CRC 1639/1 NuMeriQS – 511713970. Numerical simulations were undertaken on Barkla (though simulations with EFA could have easily been run on a laptop), part of the High Performance Computing facilities at the University of Liverpool, UK.
@@ C.1 Why heat bath cannot be generalised
The heat bath algorithm [PhysRevD.21.2308, Gattringer:2010zz] is an efficient way to sample new link variables locally in pure gauge theory. It effectively removes the potential rejection in standard Metropolis Hastings sampling. The heat bath uses the property that the gauge action $S_{G}$ is quadratic and diagonal in the plaquette $U_{\mu\nu}(n)$
@@ C.2 Derivation of the approximative FA
The optimal choice of the HMC Hamiltonian as by theorem 1 is impossible because $M_{G}=\Omega_{G}\cdot\operatorname{diag}\left(\omega_{G}^{2}\right)\cdot\Omega_{G}^{\dagger}$ is not invertible. Instead, for this work we use
$\mathcal{H}$ $=\frac{1}{2}p^{\mathsf{T}}\tilde{M}_{G}^{-1}p+S_{G}(x)$ (68)
with the regularised inverse
$\tilde{M}_{G}^{-1}$ $\coloneqq\Omega_{G}\cdot\operatorname{diag}\left(\tilde{\omega}_{G}^{-2}\right)\cdot\Omega_{G}^{\dagger}\,,$ (69)
$\tilde{\omega}_{G}^{-1}$$\coloneqq\begin{cases}\omega_{G}^{-1}&\text{if }\omega_{G}\neq 0\,,\\ \varepsilon_{G}&\text{if }\omega_{G}=0\,,\end{cases}$ (70)
where $\varepsilon_{G}=10^{-3}$ was chosen small without extensive tuning.
@@ C.3 Numerical verification of the approximate FA
For large $\beta$ the weak coupling expansion [Creutz:1983njd]
$\left\langle P(\beta)\right\rangle$ $=1-2/\beta+\operatorname{\mathcal{O}}\left(\beta^{-2}\right)$ (73)
becomes applicable. As expected, we find excellent agreement between (a) both numerical methods and the strong coupling expansion for $\beta\lesssim 4$, (b) both numerical methods and the weak coupling expansion for $\beta\gtrsim 14$, and (c) the two numerical methods with each other for the entire tested parameter range.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110022224444666688881010101012121212141414141616161618181818Refer to captionFA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=\nicefrac{{1}}{{\sqrt{\beta}}}$strong coupling expansionweak coupling expansion
@@
$\beta$
$10^{-6}$$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$00111122223333444455556666Refer to captionFA, $T=\nicefrac{{\pi}}{{2}}$no FA, $T=\nicefrac{{1}}{{\sqrt{\beta}}}$
@@
$\beta$
Figure 5: Plaquette results in SU$(3)$ pure gauge theory on a $10^{4}$ lattice from HMC simulations checked against the strong coupling expansion [PhysRevD.11.2104, PhysRevD.19.2514] using eq. (72) (the 16th order expansion was converted to Padé approximation form in order to maximise the range of validity) and the weak coupling expansion [Creutz:1983njd] using eq. (73). Left: plaquette expectation value $\left\langle P(\beta)\right\rangle$; Right: absolute value of the deviation of $\left\langle P(\beta)\right\rangle$ from eq. (72).
@@ C.4 Alternative approach sampling the plaquette
From section C.1 we learn that we have to sample in the algebra, as is done in classical lattice QCD HMC simulations for exactly this reason. One alternative option to sampling the links as in section C.2 is to still sample the plaquette like heat bath does. This means that not all the links would be updated at the same time. Instead, for every update step, one would have to choose a subset of links to update while keeping the rest fixed. Such a subset can be extensive in the volume, so global updates would still be possible. Expand the plaquette $U_{\mu\nu}(n)=\exp\left(\operatorname{i}X_{n,\mu\nu}\right)$ around unity to obtain $S_{G}\sim\beta X^{2}$. Now sample momenta according to this harmonic part and then evolve the EOM using the force from the full action. Finally update the links within the chosen subset using the changed plaquettes and multiplying be the inverse sum of fixed staples (same as heat bath).